OpenStudy (anonymous):

How do I simplify 6/√3+2? How do I rationalize the denominator to remove the radical? Do I need to do FOIL on the bottom, and do I need to multiply the denominator by the top?

OpenStudy (anonymous):

\[\huge \frac{6}{\sqrt 3} +2\] is this your problem?

OpenStudy (anonymous):

or \[\huge \frac{6}{\sqrt 3 +2} \] is your problem?

OpenStudy (anonymous):

@shatteredskittlez

OpenStudy (anonymous):

If it is \[\huge \frac{6}{\sqrt 3 +2}\] then multiply Nr. and Dr. by the conjugate of \[\huge \sqrt3 +2 \] which is \[\huge \sqrt 3-2\] and then simplify it.

OpenStudy (anonymous):

sorry my mom just figured out the answer, and this is what she did! XD thx though!